What is the density of a sample of argon gas at 70 ∘C and 785 mmHg ?

What is the density of a sample of argon gas at 70  and 785  ?

1.47 g/L
14.66 g/L
7.18 g/L

1114.19 g/L

Aluminum reacts with chlorine gas to form aluminum chloride.

2Al(s)+3Cl2(g)→2AlCl3(s)

What minimum volume of chlorine gas (at 298 Kand 279 mmHg ) is required to completely react with 8.55 g of aluminum?

What minimum volume of chlorine gas (at 298  and 279  ) is required to completely react with 8.55  of aluminum?

317 L
31.7 mL
31.7 L
0.317 L
Answer #1

1)

P= 785.0 mm Hg

= (785.0/760) atm

= 1.0329 atm

T= 70.0 oC

= (70.0+273) K

= 343 K

Molar mass of Ar = 39.95 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

1.0328947368421053 atm *39.95 g/mol = density * 0.08206 atm.L/mol.K *343.0 K

density = 1.47 g/L

Answer: 1.47 g/L

2)

Molar mass of Al = 26.98 g/mol

mass of Al = 8.55 g

mol of Al = (mass)/(molar mass)

= 8.55/26.98

= 0.3169 mol

According to balanced equation

mol of Cl2 formed = (3/2)* moles of Al

= (3/2)*0.3169

= 0.4754 mol

Given:

P = 279.0 mm Hg

= (279.0/760) atm

= 0.3671 atm

n = 0.4754 mol

T = 298.0 K

use:

P * V = n*R*T

0.3671 atm * V = 0.4754 mol* 0.08206 atm.L/mol.K * 298 K

V = 31.7 L

Answer: 31.7 L